# Towards Generalizable Scene Change Detection

### v3

@@ Abstract
* Corresponding author
@@ 3.4 Geometric-Semantic Mask Matching
Semantic Similarity Matching. Although GIM provides reasonable masks for potential object-level changes, a set of unchanged masks is included due to noise in the initial pseudo-masks. To address this issue, we perform semantic verification of the overlapping regions by calculating the cosine similarity between the mask embeddings of the bi-temporal images. Specifically, for each overlapping mask $\overline{m}_{o}$, we extract the corresponding mask embeddings $\mathcal{M}_{l,o}^{t_{0}}$ and $\mathcal{M}_{l,o}^{t_{1}}$ from the $l$-th image embeddings $\mathbf{E}_{l}$ at times $t_{0}$ and $t_{1}$, respectively. We then compute a change confidence score using the cosine similarity $c(\mathcal{M}_{l,o}^{t_{0}},\mathcal{M}_{l,o}^{t_{1}})$ which allows further refinement of the masks selected by GIM and generates the final change mask $\mathbf{Y}_{\text{pred}}$. Through a layer-wise analysis, we empirically observe that semantic differences are more pronounced in the last layer compared to the initial and intermediate layers, thereby utilizing final image embeddings for our SSM process [^1] Please refer to the supplementary material for the analysis..
@@ 5.1 Comparative Studies
Qualitative Comparison. As illustrated in Fig. 1, GeSCF adeptly adapts to various unseen images, accurately delineating changed objects—significantly outperforming the current state-of-the-art model. Moreover, as a zero-shot framework, GeSCF does not learn annotation biases in a dataset (see Fig. 5). For instance, GeSCF generates a more accurate and interpretable mask for the object and identifies unannotated changes, such as a sign, a basket, and a car. These results suggest that GeSCF’s outputs are more reasonable than baselines, even though its quantitative results are lower for some samples in the in-domain settings.
@@ 5.2 Ablation Studies
To understand the effectiveness of our contributions, we conducted comprehensive ablation studies across multiple datasets, including the standard SCD datasets and ChangeVPR (see Tab. 3). We utilized baselines trained on TSUNAMI [42], which demonstrated superior average performance over others.

### v4

@@ Abstract
Code and dataset are available here.* Corresponding author
@@ 3.4 Geometric-Semantic Mask Matching
Semantic Similarity Matching. Although GIM provides reasonable masks for potential object-level changes, a set of unchanged regions is included due to noise in the initial pseudo-masks. To address this issue, we perform semantic verification of the overlapping regions by calculating the cosine similarity between the mask embeddings of the bi-temporal images. Specifically, for each overlapping mask $\overline{m}_{o}$, we extract the corresponding mask embeddings $\mathcal{M}_{l,o}^{t_{0}}$ and $\mathcal{M}_{l,o}^{t_{1}}$ from the $l$-th image embeddings $\mathbf{E}_{l}$ at times $t_{0}$ and $t_{1}$, respectively. We then compute a change confidence score using the cosine similarity $c(\mathcal{M}_{l,o}^{t_{0}},\mathcal{M}_{l,o}^{t_{1}})$ which allows further refinement of the masks selected by GIM and generates the final change mask $\mathbf{Y}_{\text{pred}}$. Through a layer-wise analysis, we empirically observe that semantic differences are more pronounced in the last layer compared to the initial and intermediate layers, thereby utilizing final image embeddings for our SSM process [^1] Please refer to the supplementary material for the analysis..
@@ 5.1 Comparative Studies
Qualitative Comparison. As illustrated in Fig. 1, GeSCF adeptly adapts to various unseen images, accurately delineating changed objects—significantly outperforming the current state-of-the-art model. Moreover, as a zero-shot framework, GeSCF does not learn annotation biases in a dataset (see Fig. 5). For instance, GeSCF generates a more accurate and interpretable mask for the object and identifies unannotated changes—displaying resilience to erroneous GTs. These results suggest that GeSCF’s outputs are more reasonable than baselines, even though its quantitative results are lower for some samples in the in-domain settings.
@@ 5.2 Ablation Studies
To understand the effectiveness of our contributions, we conducted comprehensive ablation studies across three standard SCD datasets and ChangeVPR (see Tab. 3). We utilized baselines trained on TSUNAMI [42], which demonstrated superior average performance over others.
@@ 6 Conclusion
Acknowledgement. This research was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No.RS-2022-II220926, Development of Self-directed Visual Intelligence Technology Based on Problem Hypothesis and Self-supervised Methods; No.RS-2022-II220907, Development of AI Bots Collaboration Platform and Self-organizing AI; No.2019-0-01842, Artificial Intelligence Graduate School Program (GIST)), National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.NRF-2022R1C1C1009989), and ’Project for Science and Technology Opens the Future of the Region’ program through INNOPOLIS FOUNDATION founded by Ministry of Science and ICT (No.2022-DD-UP-0312-02-101).